The first overtone frequency of a closed organ pipe P1 is equal to the fundamental frequency of an open organ pipe P2. If the length of the pipe P1 is 30 cm, what will be the length of P2?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Let the length will be l. Here given that V = 340 m/s and n = 20 Hz Here λ/2 = l ⇒ λ = 2l We know V = nλ⇒ l = V/n = 340/2 * 20 = 34/4 = 8.5 cm (for maximum wavelength, the frequency is minimum).